# library.bib

### v1

@@ Abstract
This work presents the first large-scale neutral benchmark experiment focused on single-event, right-censored, low-dimensional survival data. Benchmark experiments are essential in methodological research to scientifically compare new and existing model classes through proper empirical evaluation. Existing benchmarks in the survival literature are often narrow in scope, focusing, for example, on high-dimensional data. Additionally, they may lack appropriate tuning or evaluation procedures, or are qualitative reviews, rather than quantitative comparisons. This comprehensive study aims to fill the gap by neutrally evaluating a broad range of methods and providing generalizable conclusions. We benchmark 18 models, ranging from classical statistical approaches to many common machine learning methods, on 32 publicly available datasets. The benchmark tunes for both a discrimination measure and a proper scoring rule to assess performance in different settings. Evaluating on 8 survival metrics, we assess discrimination, calibration, and overall predictive performance of the tested models. Using discrimination measures, we find that no method significantly outperforms the Cox model. However, (tuned) Accelerated Failure Time models were able to achieve significantly better results with respect to overall predictive performance as measured by the right-censored log-likelihood. Machine learning methods that performed comparably well include Oblique Random Survival Forests under discrimination, and Cox-based likelihood-boosting under overall predictive performance. We conclude that for predictive purposes in the standard survival analysis setting of low-dimensional, right-censored data, the Cox Proportional Hazards model remains a simple and robust method, sufficient for practitioners.
@@ 1 Introduction
Survival analysis is an important branch of statistics for data where the outcome is the time until an event of interest occurs. Such data often exhibits incomplete information about the outcome, for example due to censoring. Traditionally applied in medical research to estimate how patient survival relates to features, it has been applied in a broad range of applications across various domains. By effectively incorporating information from both completed and ongoing (censored) cases, survival analysis can yield accurate and informative predictions. This capability is invaluable in fields such as medicine, finance, and in different industrial sectors, where risk prediction is an important decision making component. Many methods have been introduced in this field, from the Cox Proportional Hazards (CPH) [cox1975] and Accelerated Failure Time (AFT) model [kalbfleisch2011] to tree-based methods including Random Survival Forests (RSFs) [ishwaran2008] and Gradient Boosting Machines (GBMs) [friedman1999, chen2016xgboost] as well as many others.
Throughout this paper, we consider only the right-censored survival setting. This restriction is due to a lack of comprehensive support for other settings (e.g. left censoring or competing risks) in available methods and especially their software implementation. In a nutshell, right-censoring occurs when some subjects did not experience the event of interest either because of drop-out or end of study (both assumed to be unrelated to the event of interest here). Formally, let $Y_{i}\sim F_{Y};Y_{i}>0;i=1,\ldots,n$ be a random variable representing the event times and $C_{i}\sim F_{C};C_{i}>0$ the censoring time. In right censored data, we do not observe realizations of $Y_{i}$ but rather of the tuple $(T_{i}=\min(Y_{i},C_{i}),D_{i}=\mathbb{I}(Y_{i}\leq C_{i}))$. The goal of survival analysis is to obtain estimates/predictions for the distribution $F_{Y}$, or quantities derived from it, e.g. $\mathbb{E}_{Y}(Y)$, based on realizations of $(T_{i},D_{i})$; see Section 3.3 for definition of prediction types. The observed data is given by tuples $(t_{i},d_{i},\mathbf{x}_{i}),\ i=1,\ldots,n$, where $t_{i}$ is the observed outcome time (either event or censoring time, whichever occurred first), $d_{i}$ is the status indicator (0 if observation is censored, 1 if the event of interest was observed) and $\mathbf{x}_{i}$ is the feature vector.
@@
Contributions: This paper introduces what we believe to be the first large-scale comparison study for single-event, right-censored survival data due to the number and range of datasets (32), models (18), tuning measures (2), and evaluation measures (8) included. We benchmark survival techniques in the the low-dimensional setting, which represents a type of data that practitioners often encounter. While ex ante one might suspect that machine learning methods will not perform well in low-dimensional settings, there are no established rules for sample sizes required by machine learning methods in the context of survival analysis. The size of our study and the inclusion of many diverse, and hopefully representative, data sets should ensure that its results can be generalized within other right-censored, low-dimensional settings [herrmann2020]. Furthermore, we run a “neutral” benchmark in accordance with the guidelines laid out by Boulesteix et al. [boulesteix2013] which we detail in Section 3.1. Based on our review of the literature (see Section 2), there is no other study: a) with a comparable number of datasets or methods; b) that compared methods after sufficient tuning for both discrimination and overall predictive ability (as measured by scoring rules); or c) that neutrally compares methods. Finally, our collection of datasets is available as an OpenML benchmark suite [openml2013] and our hyperparameter search spaces will be available in a forthcoming release of mlr3tuningspaces [pkgmlr3tuningspaces].
@@ 2 Literature Review
The experiments described in this paper provide a comparison of both classical and machine learning (ML) survival models in a low-dimensional setting. We use these broad terms for model classes in analogy to the taxonomy provided by [wang2019], where the term “classical” refers to semi- and fully parametric methods such as the CPH and AFT models or derivatives thereof, including penalized variants. “ML” here refers to non-linear and non-parametric methods ranging from tree-based methods including RSFs, boosting approaches such as GBMs or likelihood boosting (CoxBoost), to Survival Support Vector Machines (SSVMs), Artificial Neural Networks (ANNs) and Deep Learning (DL) methods. Although there is no objective and generally accepted definition to determine whether a dataset is “high-dimensional”, we colloquially define it to refer to scenarios where the number of features exceeds the number of observations ($p>n$).
Historically, surveys, reviews, and analytical comparisons of survival models can be grouped into: i) empirical comparisons of models with limited scope; ii) qualitative surveys without benchmark experiments.
@@
We provide a short overview of the literature for comparisons of survival models, with an extended review available in Appendix A. Papers that empirically compare survival models are further separated into studies that: i) compare ‘classical’ models only; ii) compare multiple ML and classical model classes; iii) compare one novel model (or class) to one or more baseline models; iv) exclusively focus on high-dimensional data.
@@
Comparisons of Classical Models often compare CPH and AFT models, including [moghimi-dehkordi2008, georgousopoulou2015, zare2015, habibi2018] showing both methods yielding similar hazard ratios but without evaluation metrics on independent test data, relying on graphical procedures to draw conclusions; [dirick2017] additionally compare flexible Cox models including splines using time-dependent AUC, MSE, and MAE, finding that the CPH with penalized splines outperformed the other models.
@@
Comparisons of ML and Classical Models The experiments carried out in this paper belong in this category. Only two prior experiments could be found that neutrally benchmarked more than one ML model class on low-dimensional data. Kattan (2003) [kattan2003] benchmarked tree-based models, ANNs and CPH with Harrell’s C-index across three datasets. The models are compared for significant differences using 50 times repeated nested cross-validation but do not clarify their tuning procedure. Boxplots across all replications indicate that no machine learning model outperformed the CPH. The authors note the small number of datasets used for comparison as their primary limitation. Zhang et al. (2021) [zhang2021survbenchmark] compare classical and ML methods, taking into account feasibility and computational efficiency for various tasks in the biomedical field. Methods are evaluated on six clinical and 16 omics datasets using 11 metrics, including time-dependent AUC, Brier score and multiple variations of the C-index. Methods were applied with specific hyperparameter settings without tuning, which limits the generalizability of their results.
@@
Comparisons of a Novel Model Class include [luxhoj1997, ohno-machado1997] benchmarking newly developed ANNs against CPH and [goli2016a] comparing new SVMs against CPH, where neither found significant differences. [jaeger2024] compare a novel implementation of oblique RSFs (“aorsf”) to the previous implementation, as well as other RSFs, GBMs, penalized CPH, and ANNs. Using Harrell’s C and ISBS, they find aorsf to outperforming GBMs and penalized CPH but with only minimal tuning applied for some models.
@@
Comparisons on High-Dimensional Data gained popularity in the survival literature, with many recent studies focusing on the area of multi-omics data [zhao2024tutorialsurvival]. [herrmann2020] perform a large-scale benchmark including penalized regression, GBMs, and RSFs, evaluating on Harrell’s C and ISBS but not finding any significant differences. [spooner2020] compare a similar group of models evaluating on Harrell’s C only, finding few significant differences. [wissel2023systematic] compare DL methods, RSFs and CPH, evaluating on Antolini’s C-index and ISBS and noting a lack of noise-resistance of all models.
@@ 3.1 Study Design
The experiments in this study are designed to assess the status quo of survival models, including both classical and machine learning approaches. In order to achieve this objective, this study aims to be a “neutral comparison study” [boulesteix2013]. Following the guidelines put forward by Boulesteix et al.:
@@
The authors are neutral: At least one representative of all methods compared in this experiment was contacted, and hyperparameter configurations were discussed with all who responded. Every party with a personal interest in the results was provided equal opportunity to influence the experiment and thereby making sure that there was no bias involved in model configuration. We are grateful for the maintainers’ time supporting this effort.
Model, performance measures and data are chosen in a rational way: The study is designed to assess the status quo, which excludes models and measures that have been published without proper peer-review. We use one primary measure each for discrimination and overall predictive ability, while additional measures are reported for comparison purposes. In order to be in line with common procedures and to allow for general comparability, we also assess models by measures even if these are known to be flawed, e.g., increasing bias of Harrell’s C for increasing censoring percentages (see, for example, [schmid2012]). The inclusion criteria for datasets were as follows: We use real-world datasets that include at least two features, a right-censoring indicator and a survival time, have at least 100 observed events, and which do not qualify as high-dimensional, i.e. have fewer features than observations. We explicitly exclude datasets with competing risk endpoints, recurrent events, or other non-standard settings such as truncation or left-censoring. No quota was specified regarding censoring proportions in the datasets.
Implementation, Reproducibility, and Accessibility Experiments were conducted on R 4.2.2 on the Beartooth Computing Environment [beartooth]. All code required to run the experiments and generate the results is available in a public GitHub repository [^1] https://github.com/slds-lmu/paper_2023_survival_benchmark licensed under GPL-3. Further details on software used are available in Appendix E. For additional reproducibility, our hyperparameter search spaces will be published with the next release of mlr3tuningspaces [pkgmlr3tuningspaces] and our datasets will be available as an OpenML benchmark suite [openml2013].
@@ 3.2 Models and Configurations
The models compared in this experiment were chosen by identifying commonly used models with readily available implementations: i) Kaplan-Meier (KM) [kaplanmeier1958]; ii) Nelson-Aalen (NA) [aalen1978]; iii) Akritas Estimator (AK) [akritas1994]; iv) Cox PH (CPH) [cox1975]; v) CV Regularized CPH (GLMN) [simon2011]; vi) Penalized (Pen) [goeman2010l1penalized]; vii) Parametric AFT (AFT)  [kalbfleisch2011]; viii) Flexible Splines (Flex) [roystonparmar2002]; ix) Random Survival Forest (RFSRC) [ishwaran2008]; x) Random Survival Forest (RAN) [ishwaran2008, pkgranger]; xi) Conditional Inference Forest (CIF) [hothorn2006]; xii) Oblique Random Survival Forest (ORSF) [jaeger2024]; xiii) Relative Risk Tree (RRT) [breiman1984]; xiv) Model-Based Boosting (MBST) [buhlmann2003boostingl2]; xv) CoxBoost (CoxB) [binder2008allowing]; xvi) XGBoost with Cox objective (XGBCox) [chen2016xgboost]; xvii) XGBoost with AFT objective (XGBAFT) [barnwal2022]; xviii) SSVM-Hybrid (SSVM) [vanbelle2011supportvector] .
The full table of all models including respective software packages and versions is given in Appendix C. In our selection we focused on well-established models with robust implementations, provided as well maintained packages or wrapper functions within benchmarking software. This excludes some recently proposed DL based methods like DeepSurv [pkgdeepsurv] and DeepHit [pkgdeephit], which have higher computational complexity, require intensive tuning, and in initial experiments could not be evaluated reliably within our benchmark suite. The KM and NA estimators are used as non-parametric baselines while AK acts as a more flexible baseline as it estimates a conditional survival function without assuming uninformative censoring. An additional table in Appendix D lists the models’ hyperparameter and pre-processing configurations. Some models could be considered equivalent as they implement the same method, e.g. RAN and RFSRC both implement Random Survival Forests, yet for the purposes of this benchmark we consider the implementation a part of the model comparison.

### v2

# arXiv:2406.04098v2[stat.ML] 28 Feb 2026

@@ Abstract
Motivation: This work presents the first large-scale neutral benchmark experiment focused on single-event, right-censored, low-dimensional survival data. Benchmark experiments are essential in methodological research to scientifically compare new and existing model classes through proper empirical evaluation. Existing benchmarks in the survival literature are smaller in scale regarding the number of used datasets and extent of empirical evaluation. They often lack appropriate tuning or evaluation procedures, while other comparison studies focus on qualitative reviews rather than quantitative comparisons. This comprehensive study aims to fill the gap by neutrally evaluating a broad range of methods and providing generalizable guidelines for practitioners.
Results: We benchmark 19 models, ranging from classical statistical approaches to many common machine learning methods, on 34 publicly available datasets. The benchmark tunes models using both a discrimination measure (Harrell’s C-index) and a scoring rule (Integrated Survival Brier Score), and evaluates them across six metrics covering discrimination, calibration, and overall predictive performance. Despite superior average ranks in overall predictive performance from individual learners like oblique random survival forests and likelihood-based boosting, and better discrimination rankings from multiple boosting- and tree-based methods as well as parametric survival models, no method significantly outperforms the commonly used Cox proportional hazards model for either tuning measure.
We conclude that for predictive purposes in the standard survival analysis setting of low-dimensional, right-censored data, the Cox Proportional Hazards model remains a simple and robust method, sufficient for most practitioners.
Availability and Implementation: All code, data, and results are publicly available on GitHub https://github.com/slds-lmu/paper_2023_survival_benchmark
Contact: burk@leibniz-bips.de
Keywords: survival analysis, risk prediction, machine learning, neutral comparison, benchmark
@@ 1 Introduction
Survival analysis is an important branch of statistics for data where the outcome is the time until an event of interest occurs. Such data often exhibits incomplete information about the outcome, for example due to censoring. Traditionally applied in medical research to estimate how patient survival relates to clinical or biological characteristics (features), it has been applied in a broad range of applications across various domains. By effectively incorporating information from both completed and ongoing (censored) cases, survival analysis can yield accurate and informative predictions. This capability is invaluable in fields such as medicine, finance, and in different industrial sectors, where risk prediction under censoring is an important decision making component. Many methods have been introduced in this field, from the Cox Proportional Hazards (CPH) (Cox, 1975, 1972) and Accelerated Failure Time (AFT) models (Kalbfleisch and Prentice, 2011) to tree-based methods including Random Survival Forests (RSFs) (Ishwaran et al., 2008) and Gradient Boosting Machines (GBMs) (Friedman, 1999; Chen and Guestrin, 2016; Barnwal et al., 2022) as well as many others. Throughout this paper, we consider only the right-censored, single-event survival setting, which is very common in practical applications. More complex settings with competing risks and different types of censoring and truncation are becoming more relevant, but are currently not broadly supported by machine learning algorithms, particularly their respective implementations. In a nutshell, right-censoring occurs when some subjects did not experience the event of interest, either due to drop-out or the end of the study (both assumed to be unrelated to the event of interest here). Formally, let $Y_{i}\sim F_{Y};Y_{i}>0;i=1,\ldots,n$ be random variables representing the event times and $C_{i}\sim F_{C};\ C_{i}>0$ the censoring time. In right-censored data, we do not observe realizations of $Y_{i}$ but rather of the tuple $(T_{i}=\min(Y_{i},C_{i}),D_{i}=\mathbb{I}(Y_{i}\leq C_{i}))$. The goal of survival analysis is to estimate the distribution $F_{Y}$, or quantities derived from it, e.g., the expected survival time $\mathbb{E}_{Y}(Y)$, based on realizations of $(T_{i},D_{i})$ and features $\mathbf{x}_{i}$; see Section 3.3 for definition of prediction types. The observed data is given by tuples $(t_{i},d_{i},\mathbf{x}_{i}),\ i=1,\ldots,n$, where $t_{i}$ is the observed outcome time (either event or censoring time, whichever occurred first), $d_{i}$ is the status indicator (0 if observation is censored, 1 if the event of interest was observed) and $\mathbf{x}_{i}$ is the feature vector.
@@
Contributions: This paper introduces the first neutral, large-scale comparison study for single-event, right-censored survival data with a large number of datasets (34), models (19), tuning measures (2), and evaluation measures (6) included. We benchmark survival techniques in the low-dimensional setting, which represents a type of data that practitioners often encounter. The scale of our study, encompassing diverse datasets (see Table 2) and an extensive evaluation procedure using multiple metrics, provides a robust benchmark for comparing model performance in right-censored, low-dimensional settings (Herrmann et al., 2021). Furthermore, we run a “neutral” benchmark in accordance with the guidelines laid out by Boulesteix et al. (2013) which we further outline in Section˜3.1. Based on our review of the literature (see Section˜2), there is no other study: a) with a comparable number of datasets or methods; b) that compared methods after sufficient tuning for both discrimination and for overall predictive ability as measured by the integrated survival Brier score (ISBS); or c) that neutrally compares methods. Finally, our experiment code is available on GitHub and our collection of datasets is available as an OpenML benchmark suite (Vanschoren et al., 2013) and our hyperparameter search spaces will be available in a forthcoming release of mlr3tuningspaces (Becker, 2024).
@@ 2 Literature Review
The experiments described in this paper provide a comparison of both classical and machine learning (ML) survival models in a low-dimensional setting. We use these broad terms for model classes in analogy to the taxonomy provided by Wang et al. (2019), where the term “classical” refers to semi- and fully parametric methods such as the CPH and AFT models or derivatives thereof, including penalized variants. “ML” here refers to non-linear and non-parametric methods ranging from tree-based methods including RSFs, boosting approaches such as GBMs or likelihood-based boosting (CoxBoost), to Survival Support Vector Machines (SSVMs), Artificial Neural Networks (ANNs) and Deep Learning (DL) methods. Although there is no objective and generally accepted definition to determine whether a dataset is “high-dimensional”, we colloquially define it to refer to scenarios where the number of features exceeds the number of observations ($p>>n$).
@@
Historically, surveys, reviews, and analytical comparisons of survival models can be grouped into: 1) empirical comparisons of models with limited scope; 2) qualitative surveys without benchmark experiments. We provide a short overview of the literature for comparisons of survival models, with an extended review available in Appendix A. Papers that empirically compare survival models (1) are further separated into studies that: (i) compare ‘classical’ models only; (ii) compare multiple ML and classical model classes; (iii) compare one novel model (or class) to one or more baseline models; (iv) exclusively focus on high-dimensional data.
@@
Comparisons of Classical Models often compare CPH and AFT models, including (Moghimi-Dehkordi et al., 2008; Georgousopoulou et al., 2015; Zare et al., 2015; Habibi et al., 2018) showing both methods yielding similar hazard ratios but without evaluation metrics on independent test data, relying on graphical procedures to draw conclusions; (Dirick et al., 2017) additionally compare flexible Cox models including splines using time-dependent area under the Receiver Operating Characteristic curve (AUC), mean squared error, and mean absolute error, finding that the CPH with penalized splines outperform the other models.
@@
Comparisons of ML and Classical Models The experiments carried out in this paper belong in this category. Only two prior experiments could be found that neutrally benchmarked more than one ML model class on low-dimensional data. Kattan (2003) benchmarked tree-based models, ANNs and CPH with Harrell’s C-index across three datasets. The models are compared for significant differences using 50 times repeated nested cross-validation, but the authors do not clarify their tuning procedure. Boxplots across all replications indicate that no machine learning model outperformed the CPH. The authors note the small number of datasets used for comparison as their primary limitation. Zhang et al. (2021) compare classical and ML methods, taking into account feasibility and computational efficiency for various tasks in the biomedical field. Methods are evaluated on six clinical and 16 omics datasets using 11 metrics, including time-dependent AUC, ISBS and multiple variations of the C-index. Critically, methods were applied with implementation-specific default hyperparameter settings without tuning, severely limiting the generalizability of their results.
@@
Comparisons of a Novel Model Class include Luxhoj and Shyur (1997); Ohno-Machado (1997) benchmarking newly developed ANNs against CPH and Goli et al. (2016) comparing modern Support Vector Regression variants against CPH, where neither found significant differences. Jaeger et al. (2024) compare a novel implementation of oblique RSFs (ORSF) to the previous implementation, as well as other RSFs, GBMs, penalized CPH, and ANNs. Using Harrell’s C and ISBS, they find that ORSFs outperform GBMs and penalized CPH, but with only minimal tuning applied for some models.
@@
Comparisons on High-Dimensional Data have gained popularity in the survival literature, with many recent studies focusing on the area of multi-omics data (Zhao et al., 2024). Herrmann et al. (2021) perform a large-scale benchmark including penalized regression, GBMs, and RSFs, evaluating on Harrell’s C and ISBS without finding any significant differences between the different model classes . Spooner et al. (2020) compare a similar group of models evaluating on Harrell’s C only, finding few significant differences. Wissel et al. (2023) compare DL methods, RSFs and CPH, evaluating on Antolini’s C-index and ISBS with a focus on noise-resistance, noting a lack thereof for all models.
@@ 3.1 Study Design
The experiments in this study are designed to assess the status quo of survival models, including both classical and machine learning approaches. In order to achieve this objective, this study aims to be a “neutral comparison study” (Boulesteix et al., 2013). Following the guidelines put forward by Boulesteix et al., such studies:
@@
Fair and neutral setup: At least one representative of all methods compared in this experiment was contacted, and hyperparameter configurations were discussed with all who responded. Every maintainer of the used software packages implementing the evaluated methods was given equal opportunity to influence the experiment, ensuring there was no bias in model configuration. We are grateful for the maintainers’ time supporting this effort.
Model, performance measures, and data are chosen in a rational way: The study is designed to assess the status quo, which excludes models and measures that have been published without peer-review. We focus on a single primary metric to evaluate discrimination and another for overall predictive performance, while additional measures are reported for comparison purposes. To align with common practice and to allow for general comparability, we also assess models by measures even if these are known to be flawed, e.g., increasing bias of Harrell’s C-index for increasing censoring percentages (see, for example, Schmid and Potapov (2012)). The inclusion criteria for datasets were as follows: We use real-world datasets that include at least two features, a right-censoring indicator with a survival time, have at least 100 observed events, and do not qualify as high-dimensional, i.e., have fewer features than observations. We exclude datasets with competing risk endpoints, recurrent events, or other non-standard settings such as left-censoring. No quota was specified regarding censoring proportions in the datasets.
Implementation, Reproducibility, and Accessibility Experiments were conducted on R 4.4.3 on the Beartooth Computing Environment (Beartooth Computing Environment, x86_64 cluster, 2024). All code required to run the experiments and generate the results is available in a public GitHub repository [^1] https://github.com/slds-lmu/paper_2023_survival_benchmark licensed under GPL-3. Further details on software used are available in Appendix D. For additional reproducibility, our hyperparameter search spaces will be published with an upcoming release of mlr3tuningspaces (Becker, 2024) and our datasets will be available as an OpenML benchmark suite (Vanschoren et al., 2013) while already being available from GitHub.
@@ 3.2 Models and Configurations
The algorithms compared in this experiment were chosen by identifying commonly used models with readily available implementations: (1) Kaplan-Meier (KM) (Kaplan and Meier, 1958); (2) Nelson-Aalen (NEL) (Aalen, 1978); (3) Akritas Estimator (AK) (Akritas, 1994); (4) Cox PH (CPH) (Cox, 1975); (5) Regularized CPH with cross-validation (GLMN) (Simon et al., 2011); (6) Penalized (Pen) (Goeman, 2010); (7) Parametric AFT (AFT) (Kalbfleisch and Prentice, 2011); (8) Flexible Splines (Flex) (Royston and Parmar, 2002); (9) Random Survival Forest (RFSRC) (Ishwaran et al., 2008); (10) Random Survival Forest (RAN) (Wright and Ziegler, 2017); (11) Conditional Inference Forest (CIF) (Hothorn et al., 2006); (12) Oblique Random Survival Forest (ORSF) (Jaeger et al., 2024); (13) Relative Risk Tree (RRT) (Breiman, 1998); (14) Model-Based Boosting with Cox and AFT objective (MBSTCox and MBSTAFT) (Bühlmann and Yu, 2003); (15) CoxBoost (CoxB) (Binder and Schumacher, 2008); (16) XGBoost with Cox and AFT objective (XGBCox and XGBAFT) (Chen and Guestrin, 2016; Barnwal et al., 2022); (17) Survival-SVM (SSVM) (Van Belle et al., 2011).
The full table of all models, including respective software packages, is given in Appendix C, and the online supplement includes recorded versions of all packages involved in the experiment. In our selection, we focused on well-established models with generally robust implementations, provided as well-maintained packages or wrapper functions within benchmarking software. This excludes some recently proposed DL based methods like DeepSurv (Katzman, 2016) and DeepHit (Lee, Changhee et al., 2019), which have higher computational complexity, require intensive tuning, and in initial experiments could not be evaluated reliably within our benchmark suite due to a lack of stable implementations. The KM and NEL estimators are used as non-parametric baselines, while AK acts as a more flexible baseline as it estimates a conditional survival function without relying on independent or non-informative censoring. An additional table in Appendix C lists the associated hyperparameter and pre-processing configurations. Note that while some studied methods could be considered algorithmically equivalent, e.g., RAN and RFSRC both implement Random Survival Forests, their implementations are not, and for that reason, we opted to include such variants in the comparison. A small number of methods provide internal optimization methods for certain hyperparameters, which we used when available / possible in our global tuning procedure. This affects both XGBoost models and their early-stopping mechanism to tune the nrounds parameter and GLMN for lambda, while CoxB is exclusively tuned using its internal mechanism. For further details we refer to the experiment code on GitHub.